# Yongyang Cai  Khyati Malik   Hyeseon Shin

### v2

@@ A.5 GDP Growth Rate beyond this Century
A.6 Calibration of Total Factor Productivity
Figure A.4 shows that with our calibrated $A_{i,t}$ , the GDP per capita $y_{i,t}^{\mathrm{NoCC}}$ matches well with the projected data $y_{i,t}^{\mathrm{BDD}}$ from Burke et al. (2018) for all regions.
Figure A.4: Fitting GDP per capita under no climate impact. Lines represent GDP per capita from Burke et al. (2018); marks represent fitted GDP per capita.
A.7 Calibration of Climate Damage
Figure A.5 shows that with our calibrated climate damage coefficients, the ratios of GDP per capita between RCP 2.6 and RCP8.5 from our model, matches well with the ratios in Kahn et al. (2021), for every region.
Figure A.5: Fitting climate damage parameters. Lines represent the ratios of GDP per capita between RCP 2.6 and RCP8.5 from Kahn et al. (2021); marks represent fitted ratios.
A.8 Additional Simulation Results
A.8.1 Regional Net Emissions under Noncooperation
Figure A.6 displays the regional net emissions under the noncooperative model with the ETS and the baseline emission cap scenario. Latin America is the first to reach net emissions in 2053, followed by Russia in 2055, the US in 2056, OHI in 2057, and Eurasia in 2058. Then, net emissions are achieved by China in 2061, Japan in 2062, the MidWest in 2064, and the EU in 2065. Finally, India achieves net emissions in 2069, followed by Africa and other non-OECD Asian countries in 2071.
Figure A.6: Simulation results of regional net emissions under noncooperation with the ETS and baseline emission caps.
A.8.2 Regional Net Emissions Comparison between Different Cases
Figure A.7 compares regional net emissions between two cases under cooperation with the baseline emission caps: (i) with the ETS, (ii) without the ETS.
Figure A.7: Simulation results of regional net emissions: comparison between two cases under cooperation.
Figure A.8 compares regional net emissions between two cases under noncooperation with the baseline emission caps: (i) with the ETS, (ii) without the ETS.
Figure A.8: Simulation results of regional net emissions: comparison between two cases under noncooperation.
A.8.3 Regional MAC Comparison between Different Cases
Figure A.9 compares the regional MAC under cooperation with the baseline emission caps. We compare two cases: (i) with the ETS, (ii) without the ETS.
Figure A.9: Simulation results of regional MAC: comparison between two cases under cooperation.
@@ A.8.3 Regional MAC Comparison between Different Cases
Figure A.10: Simulation results of regional MAC: comparison between two cases under noncooperation.
A.8.4 Regional SCC Comparison between Different Cases
Figure A.11 compares the regional SCC under cooperation with the baseline emission caps. We compare two cases: (i) with the ETS, (ii) without the ETS.
Figure A.11: Simulation results of the regional SCC: comparison between two cases under cooperation.
@@ A.8.4 Regional SCC Comparison between Different Cases
Figure A.12: Simulation results of the regional SCC: comparison between two cases under noncooperation.
A.8.5 Regional MAC and SCC under Alternative Emission Caps
@@
Figure A.13: Simulation results of regional MAC: comparison across alternative emission cap scenarios.
@@
Figure A.14: Simulation results of regional SCC: comparison across alternative emission cap scenarios.
A.8.6 Sensitivity Analysis on Marginal Abatement Cost
In Figure A.15, we provide additional sensitivity analysis results on the abatement cost. Recall that, the MAC in the model is
$\mathrm{MAC}{}_{i,t}=1,000b_{2,i}\mu_{i,t}^{b_{2,i}-1}(b_{1,i}+b_{3,i}\exp(-b_{4,i}t)).$
We investigate a scenario wherein the MACs across all regions experience a variation of ±20% from the baseline MAC levels at the same emission control rates $\mu_{i,t}$. Specifically, we explore the effects of adjusting two key parameters $b_{1,i}$ and $b_{3,i}$, by ±20% from their baseline values. The result shows that the equilibrium permit price is closely linked to the level of the MACs. When the MACs are elevated across all regions, the permit price also experiences a corresponding increase with almost identical changes of the MACs, and vise versa. For example, in 2040, the baseline MAC in the US is \$556, it is changed to \$644 or \$467, and the permit price is changed from \$404 to \$492 or \$315. That is, both the MACs and the permit price increase \$88 or decreases \$89 in 2040. At the same time, the regional SCC has little changes, as the existence of the emission caps makes the future temperature have little changes and the changes in the MACs have little impact on marginal climate damages in future periods and social discount rates. This again verifies our relation between the optimal carbon tax, the MAC, and the permit price (i.e., equation (20)).
Figure A.15: Sensitivity Analysis on the MAC.

### v3

@@ A.3.5 GDP Growth Rate beyond this Century
A.4 Additional Simulation Results
A.4.1 Benchmark Model: Regional Emissions
Figure A.7 displays the regional emissions under the noncooperative model with the ETS and the baseline emission cap scenario. Russia is the first to reach net zero emissions in 2050, followed by China and Latin America in 2056, MidEast in 2057, the US and Eurasia in 2058, and the OHI in 2059. Then, net zero emissions are achieved by EU in 2061, Japan and India in 2064. Finally, Africa and non-OECD Asia achieve net zero emissions in 2070.
Figure A.7: Simulation results of regional emissions under the baseline emission caps.
A.4.2 Model Comparison of ETS Implementation
Figure A.8 compares regional emissions between two cases under noncooperation with the baseline emission caps: (i) with the ETS, (ii) without the ETS.
Figure A.8: Comparison of the ETS implementation: regional emissions.
@@ A.4.2 Model Comparison of ETS Implementation
Figure A.9: Comparison of the ETS implementation: regional MAC.
@@
Figure A.10: Comparison of the ETS implementation: regional SCC.
A.4.3 Sensitivity Analysis over Emission Caps
@@
Figure A.11: Comparison of regional MAC under different emission caps.
@@
Figure A.12: Comparison of regional SCC under different.
A.4.4 Sensitivity Analysis over Climate Damage Parameters
In Figure A.13, we compare the MAC for the noncooperative model with the ETS under alternative values of the climate damage parameters ($\pi_{1,i}$, $\pi_{2,i}$), based on climate damage projections from Kahn et al., (2021) and Nordhaus, 2010a alongside the baseline climate damage projection from Burke et al., (2018).
Figure A.13: Comparison of regional MAC under different estimates of climate damages.
Similarly, Figure A.14 displays the SCC of the noncooperative model with the ETS under different values of the climate damage parameters.
Figure A.14: Comparison of regional SCC under different estimates of climate damages.
A.4.5 Sensitivity Analysis over Abatement Cost Parameters
In Figure A.15, we compare the MAC for the noncooperative model with the ETS under alternative estimates of the emissions abatement parameters ($b_{1,i}$, $b_{2,i}$, $b_{3,i},$ $b_{4,i}$), calibrated from Ueckerdt et al., (2019) and Nordhaus, 2010a .
Figure A.15: Comparison of regional MAC under different estimates of emissions abatement cost.
Similarly, Figure A.16 displays the SCC of the noncooperative model with the ETS under different emissions abatement cost estimates.
Figure A.16: Comparison of of regional SCC under different estimates of emissions abatement cost.
A.4.6 Sensitivity over TFP Growth Rates
Figure A.17 compares key model outcomes under different TFP growth rates: the baseline TFP growth rates derived from Burke et al., (2018) and the alternative rates based on Nordhaus, 2010a .
Figure A.17: Comparison of simulation results under different TFP growth rates.